40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2645     Entry time: Sun Feb 28 16:45:05 2010
Author: rana 
Type: Summary 
Category: General 
Subject: Power ON Recovery 
  1. Turned ON the RAID above linux1.
  2. Hooked up a monitor and keyboard and then turned ON linux1.
  3. After linux1 booted, turned ON nodus - then restarted apache and elog on it using the wiki instructions.
  4. Turned on all of the control room workstatiions, tuned Pandora to Johnny Cash, started the auto package updater on Rosalba (517 packages).
  5. Started the startStrip script on op540m.
  6. turned on RAID for frames - wait for it to say 'SATA', then turn on daqctrl and then fb40m and then daqawg and then dcuepics
  7. turned on all the crates for FEs, Sorensens, Kepcos for LSC, op340m, mafalda was already on
  8. fb40m again doesn't mount the RAID again!
  9. I turned on fb40m2 and that fixes the problem. The fb40m /etc/vfstab points to 198.168.1.2, not the JetStor IP address.
  10. I plugged in the Video Switch - its power cord was disconnected.
  11. FEs still timing out saying 'no response from EPICS', but Alberto is now here.

Sun Feb 28 18:23:09 2010

Hi. This is Alberto. Its Sun Feb 28 19:23:09 2010

  1. Turned on c1dcuepics, c0daqctrl and c0daqawg. c0daqawg had a "bad"status on the daqdetail medm screen. The FEs still don't come up.
  2. Rebooted c1dcuepics and power cycled c0daqctrl and c0daqawg. The problem is still there.
  3. Turned on c1omc. Problem solved.
  4. Rebooted c1dcuepics and power cycled c0daqctrl and c0daqawg. c0daqawg now good. The FE are coming up.
  5. Plugged in the laser for ETMY's oplev
  6. Turned on the laser of ETMX's oplev from its key.

 Monday, March 1, 9:00 2010 Steve turns on PSL-REF cavity ion pump HV at 1Y1

ELOG V3.1.3-